OpenStudy (anonymous):

If (y-1)^3=8 what is the value of (y+1)^2?

OpenStudy (anonymous):

(y-1)^3=8 (y-1)=8^(1/3) (y-1)=2 y=2 so, (y+1)^2=(2+1)^2 =(3)^2=9

OpenStudy (anonymous):

thanks!